2. At what distance above the surface of the earth is the acceleration due to the earth's gravity 0.835 m/s2 if the acceleration due to gravity at the surface has magnitude 9.80m/s2?

______ m

Can you finish this problem before 12 pm central time cause around that time is when the assignment is due. Thank you for your time, have a great day!
Answer #1

here,

let the distance from the surface of earth is x

as g = go/(1 + h/R)^2

where R is the radius of earth

0.835 = 9.8/(1 + h/6371)^2

solving for h

x = 15422.2 km = 1.54 *10^7 m

the distance from the surface of earth is 1.54 *107 m
Answer #2

SOLUTION :


Let that height be h  m from the earth’s surface at which altitude, g’ = 0.835 m/s^2


g at the earth’s surface is = 9.8 m/s^2.


As we go up, value of acceleration due to gravity decreases.


Value of g’ at height h is given by the following formula:


g’ = g (1 - 2h / R) 


Where,


g’ = acceleration due to gravity at height h = 0.835  m/s^2 (given in the problem)


g = acceleration due to gravity at earth’s surface = 9.8 m/s^2


h = height at which g’ = 0.835 m/s^2. (To be determined)


R = radius of the earth = 6371 km = 6.371*10^6 m


So, as per above formula :


0.835 = 9.8( 1 - 2 h / (6.371 * 10^6) )

=> h = (1 - 0.835/9.8) *  (6.371 * 10^6) / 2 

=> h = 2 914 082.40 m = 2914.082 km

=> h = height at which g’ is 0.835 m/s^2 = 2914.082 km  approx.  (ANSWER).

answered by: Tulsiram Garg